OpenStudy (anonymous):

(SQRT5)^x=125. I know it equals but I would love a step by step on how to solve

OpenStudy (phi):

is this supposed to be \[ \left( \sqrt{5}\right)^x = 125 \]

OpenStudy (anonymous):

Yes

OpenStudy (phi):

re-write the square root as 5 to the 1/ 2 power \[ \left( 5^\frac{1}{2}\right)^x = 125 \] also, it is useful to notice that 125 = 5*5*5 = 5^3 \[ \left( 5^\frac{1}{2}\right)^x = 5^3 \]

OpenStudy (phi):

and we also want to use this rule \[ \left( a^b\right)^c = a^{b\cdot c} \]

OpenStudy (anonymous):

Why 5^1/2

OpenStudy (phi):

notice that \[ 5^\frac{1}{2} \cdot 5^\frac{1}{2} = 5^{\frac{1}{2} +\frac{1}{2} } =5^1 = 5\] by definition, if you have A*A= 5 then A is the square root of 5 in other words, \[ \sqrt{5} \text{ is the *same thing* as } 5^\frac{1}{2} \]

OpenStudy (anonymous):

Ok

OpenStudy (phi):

In general, the exponent version is easier to deal with.

OpenStudy (phi):

can you apply the rule \[ \left( a^b\right)^c = a^{b\cdot c} \] to your problem \[ \left( 5^\frac{1}{2}\right)^x = 5^3 \]?

OpenStudy (phi):

you should get \[ 5^\frac{x}{2} = 5^3 \] the bases are equal , so you can equate the exponents: \[ \frac{x}{2} = 3 \] now solve for x by multiplying both sides by 2

OpenStudy (anonymous):

6

OpenStudy (phi):

yes, x=6 is the answer. to review: replace the square root with the exponent 1 /2 also, notice you have a base 5 on the left side. Try to make the right side also a base of 5 ... in other words, expect you can write 125 as some power of 5. then equate exponents.

OpenStudy (anonymous):

Ok Thanks

OpenStudy (anonymous):

I get it now